Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
mantovani, gherardini, chiummo, lumaca, spinicelli - 15:37 Monday 31 October 2022 (57603) Print this report
ITF recovery just started

This morning I found the ITF locked since less than two hours.

I've engaged the COMMp and BS ty in full bandwidth (see previous log entry).

Then I've automatized the engagment in full bandwidth of the two signals (to be restarted and checked after the scan of the cavities performed by Fabio in the hot state).

@ 7:34 UTC scan of the cavities

The metatron has been modified

1st attempt NE and BS in full bandwidth failed at carm 0 1f

Antonino has removed the control of the IB z on the FF RF QDP I H, which acts on the BPC x dof, to evaluate the effect in the next locks, see figure 1.

We started to unlock at CARM 0 3f, thus :

we reverted the action on the IB z control

we noticed that since the 29 the sidebands were higher, see fig 2. We called Diana to restore the old configuration and she noticed that the power sent to the WI das was the different wrt the one expected.

Even if the recycling was better we decided to restore the DAS value to check if the ITF could lock and to take this information for the next DAS tuning shift (CH?)

At DRMI it was visible that the sideband powers went back to the 29th value. We could relock at 3f and we unlocked then at 1f due to BS ty high frequency oscillation (closed in full bandwidth). Indeed the gain of BS was tuned after hours of lock at carm 0.

The engagment in full bandwidth has been reverted.

We could not lock anymore above 40mW

The green WARM parameter retuned:

Modification on WARM ALS

- refl high from 0.87 to 0.82

- refl low from 0.92 to 0.87

- lock treshold to 0.82

 

The recovery was not done. A complete recovery is needed and it would advisable to re-check the TCS parameters.

Images attached to this report
Comments to this report:
rocchi - 23:24 Monday 31 October 2022 (57612) Print this report

The slow variation of the WI DAS power is connected to a slow change in the laser head temperature, which is typical of Spring and Autumn.

A possibile solution is to decrease by 0.1 °C the chiller set point.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×